# Loc2: Interpretable Cross-View Localization via Depth-Lifted Local Feature Matching

### v2

@@ 4.5 Ablation Study
Our main ablation study focuses on two key design choices: (1) the coordinate assignment strategy (Sec. 3.2), comparing using all points within the depth range versus only the topmost points along height; and (2) the effect of ignoring scale in Procrustes alignment (reducing our formulation to the orthogonal Procrustes of Xia & Alahi (2025); Barroso-Laguna et al. (2024)). As shown in Tab. 3, explicitly selecting the topmost points does not improve localization accuracy. The side view of an object’s top does not necessarily exhibit the most consistent semantics with the aerial image. Therefore, enforcing the use of topmost points may hinder cross-view localization. For scale-awareness, excluding scale forces the model to align features strictly to metric depth predictions, which are imperfect. Allowing the model to also estimate scale from correspondences yields more accurate and robust matching. Moreover, without modeling scale, the method cannot support relative depth predictors, as orthogonal Procrustes assumes that all coordinates lie in the same scale space.
@@ 5 Conclusion
    Test on non-road facing samples on KITTI
@@
  8. H.
@@ Appendix A Additional ablation study
We conduct an additional ablation study on the VIGOR same-area validation set with unknown orientation, analyzing the effects of the number of sampled correspondences $N$, the number of aerial points, the aerial image resolution, and the use of RANSAC.
Number of sampled correspondences $N$: Our default setting uses $N=1024$ (following Xia & Alahi (2025)) during training. In inference, if RANSAC is used, we sample only 3 correspondences per RANSAC loop to compute the pose and scale. When RANSAC is not used, we use the same $N=1024$.
We evaluate two cases: 1) varying $N$ only in inference: a model trained with $N=1024$ is evaluated with $N=256,512,1024,$ and $2048$ (without RANSAC). 2) varying $N$ in both training and inference: validation results (with RANSAC) of models trained with $N=256,512,1024,$ and $2048$.
As shown in Tab. 4, the number of correspondences $N$ does not have a strong impact on performance. When varying $N$ during inference, the difference in mean localization error is less than 0.2 m for $N=512,1024,2048$. The error increases more noticeably when a smaller $N$, such as 256, is used. Similarly, the model remains robust to different $N$ during training, as long as $N$ is not too small (e.g., $N\leq 256$).
@@
Training
2048 3.91 1.76 9.25 1.73
1024 (default) 3.86 1.75 9.30 1.93
512 4.12 1.80 9.71 1.80
256 4.31 1.90 11.27 2.17
Number of aerial points: Next, we examine how the number of aerial points affects performance. Our default setting, following Xia & Alahi (2025), uses $41\times 41$ aerial points for fair comparison. As shown in Tab. 5, our model is robust to small variations in the number of aerial points. When more than $36\times 36$ points are used, the change in mean localization error is less than 0.2 m. However, when the number of points is significantly reduced (e.g., $31\times 31$), performance degrades noticeably.
@@ Appendix B Test on non-road facing samples on KITTI
Unknown
Cross-area (non-road) CCVPE 8.33 5.99 89.78 93.97
@@ Appendix C Additional Local Feature Matching Results
Failure cases: Typical failure cases are shown in Fig. 8. When there is dense vegetation (Fig. 8 (a),(b)), the matchable information across views is very limited, making precise localization extremely challenging, even for humans. Interestingly, in Fig. 8 (b), our method still correctly matches the gap between two buildings in the ground view to the corresponding gap in the aerial view. In Fig. 8 (c), the two narrow roads with trees and a zebra crossing in front appear visually similar. At the true location, a tree partially blocks the zebra crossing in the aerial view, making it harder to distinguish. As a result, our method matches the zebra crossing to an incorrect one, leading to an erroneous location estimate. Finally, in Fig. 8 (f), it is nearly impossible to determine the exact location over the water due to the lack of distinctive features.
We highlight again that this strong interpretability of our method stems from its explicit matches of local features, from which the output pose is analytically computed. Moreover, these matches are independent of the specific depth model used, as the depth information is only utilized when computing the pose from the correspondences.
@@
Appendix D Additional Results on Using Relative Depth in Inference
@@
Appendix E Training and Testing with Relative Depth
@@
Appendix F Additional Proof for Scale-Aware Procrustes Alignment
@@
Appendix G Additional Details on InfoNCE Losses
@@
Appendix H Runtime and Memory Usage
@@
Runtime: On a single H100 GPU, our method achieves 14.74 FPS without RANSAC (Fischler & Bolles, 1981) (0.03s for our model and 0.04s for the depth predictor Unik3D (Piccinelli et al., 2025)) and 0.42 FPS with RANSAC. This is faster than the previous local feature matching method FG2 (Xia & Alahi, 2025), which runs at 9.26 FPS without RANSAC and 0.32 FPS with it. Notably, when prioritizing speed, omitting RANSAC results in only a small increase in localization error, e.g., on VIGOR same-area test set with known orientation the mean error increases from 3.06 m to 3.29 m. Furthermore, one can use a faster depth predictor to further reduce inference time.

### v3

@@ 4.6 Ablation Study
Our main ablation study focuses on two key design choices: (1) the coordinate assignment strategy (Sec. 3.2), comparing using all points within the depth range versus only the topmost points along height; and (2) the effect of ignoring scale in Procrustes alignment (reducing our formulation to the orthogonal Procrustes of Xia and Alahi (2025); Barroso-Laguna et al. (2024)). As shown in Tab. 3, explicitly selecting the topmost points does not improve localization accuracy. The side view of an object’s top does not necessarily exhibit the most consistent semantics with the aerial image. Therefore, enforcing the use of topmost points may hinder cross-view localization. For scale-awareness, excluding scale forces the model to align features strictly to metric depth predictions, which are imperfect. Allowing the model to also estimate scale from correspondences yields more accurate and robust matching. Moreover, without modeling scale, the method cannot support relative depth predictors, as orthogonal Procrustes assumes that all coordinates lie in the same scale space. Additional ablation study on other hyperparameters in both inference and training is included in Appendix. Sec. A.
@@ 5 Conclusion
Reproducibility statement: Our method combines a deep learning model with a classic algorithmic component, scale-aware Procrustes alignment. The deep learning component is described in Sec. 3.2, while the derivation of the scale-aware Procrustes alignment is detailed in Sec. 3.2 with a complete proof provided in the Appendix, ensuring that the theoretical foundation of our approach is transparent and verifiable. The implementation is lightweight and relies only on standard components, with all hyperparameters and training settings fully described in Sec. 4.2. To further support reproducibility, we will release our code together with configuration files and evaluation scripts, enabling the community to reproduce all experiments and results reported in this paper.
@@
    1. A.1.
      Ablation on inference settings.
    2. A.2.
      Ablation on training settings.
@@
    Test on non-road facing samples on KITTI.
@@
  9. I.
    Discussion of standard and true digital orthophoto maps.
  10. J.
@@ Appendix A Additional Ablation Study
We conduct additional ablation studies on the VIGOR same-area validation set with unknown orientation, analyzing the effects of various hyperparameters in both inference and training.
During inference, we vary the number of sampled correspondences $N$, the number of aerial points, the resolution (meters per pixel) of the input aerial image, and the use of RANSAC in our trained model using the default settings.
During training, we investigate the influence of various hyperparameters, including the number of sampled correspondences $N$, the number of aerial points, the feature map resolution, and the temperature parameter $\tau$. Additionally, we compare our formulation (image-plane matching + scale-aware Procrustes alignment) with two homography-based formulation: (1) Replacing image-plane matching with BEV-plane matching by first transforming the ground images into a BEV, and then feed it to the same feature extractor, matcher, and Procrustes alignment. (2) Replacing the scale-aware Procrustes alignment with homography for pose estimation.
A.1 Ablation study on inference settings
Number of sampled correspondences $N$: Our default setting uses $N=1024$ (following Xia and Alahi (2025)) during inference. If RANSAC is used, we sample only 3 correspondences per RANSAC loop to compute the pose and scale. In this case, we vary $N$ only in inference: a model trained with $N=1024$ is evaluated with $N=256,512,1024,$ and $2048$ (without RANSAC).
As shown in Tab. 4, the number of correspondences $N$ does not have a strong impact on performance. When varying $N$ during inference, the difference in mean localization error is less than 0.2 m for $N=512,1024,2048$. The error increases more noticeably when a smaller $N$, such as 256, is used.
@@ A.1 Ablation study on inference settings
Number of aerial points: Next, we examine how the number of aerial points affects performance. Our default setting, following Xia and Alahi (2025), uses $41\times 41$ aerial points for fair comparison. As shown in Tab. 5, our model is robust to small variations in the number of aerial points. When more than $36\times 36$ points are used, the change in mean localization error is less than 0.2 m. However, when the number of points is significantly reduced (e.g., $31\times 31$), performance degrades noticeably.
@@
A.2 Ablation study on training settings
We first investigate the impact of several hyperparameters involved in our local feature matching during training, including the number of sampled correspondences, the number of aerial points, and the temperature parameter. After this, we vary our formulation by using homography-transformed inputs or a homography-based pose estimator.
Number of sampled correspondences $N$: Our default setting uses $N=1024$ (following Xia and Alahi (2025)) during training. In this case, we vary $N$ in both training and inference: validation results (with RANSAC) of models trained with $N=256,512,1024,$ and $2048$. As shown in Tab. 8, the model remains robust to different $N$ during training, as long as $N$ is not too small (e.g., $N\leq 256$).
Mode N Localization (m) Orientation ()
Mean ↓ Median ↓ Mean ↓ Median ↓
Training 2048 3.91 1.76 9.25 1.73
1024 (default) 3.86 1.75 9.30 1.93
512 4.12 1.80 9.71 1.80
256 4.31 1.90 11.27 2.17
Number of aerial points & Feature Map Resolution: We train additional models using $31\times 31$, $36\times 36$, $46\times 46$, $51\times 51$, $56\times 56$ aerial points. Meanwhile, to investigate the impact of feature map resolution, we also trained a model that upsamples the aerial DINO feature map from its original $45\times 45$ resolution to $180\times 180$ using two stages of bilinear interpolation followed by convolutional layers. We then sample a $41\times 41$ grid of points from this higher-resolution feature map.
As shown in Tab. 9, higher resolution of aerial points provides only a modest improvement in performance (0.16 meters reduction in mean localization error). However, such setting substantially increases memory usage and training time. Notably, Tab. 10 indicates that this marginal improvement mainly comes from RANSAC, which benefits from a larger pool of correspondences. When inferring without RANSAC, the $41\times 41$ setting even performs slightly better. By contrast, upsampling the feature maps does not lead to any measurable performance improvement.
Table 9: Ablation on different resolution configurations. Best in bold.
Number of aerial points ($N$) Mean loc. (m) Median loc. (m) Mean ori. () Median ori. ()
$31\times 31$ 4.22 1.90 9.68 1.99
$36\times 36$ 4.01 1.82 9.42 1.91
$41\times 41$ (default) 3.86 1.75 9.30 1.93
$41\times 41$ (upsampled feature) 4.55 1.94 11.46 1.96
$46\times 46$ 3.90 1.72 9.07 1.72
$51\times 51$ 3.75 1.67 8.76 1.73
$56\times 56$ 3.70 1.65 8.75 1.69
Table 10: Different resolution configurations (inference without RANSAC). Best in bold.
Number of aerial points ($N$) Mean loc. (m) Median loc. (m) Mean ori. () Median ori. ()
$41\times 41$ (default) 4.60 2.81 11.19 4.04
$51\times 51$ 4.70 2.89 11.38 4.03
$56\times 56$ 4.68 2.91 11.59 3.92
Temperature Parameter $\tau$: This parameter directly affects the matching score matrix $M^{*}$. Following Barroso-Laguna et al. (2024); Sun et al. (2021); Xia and Alahi (2025), we adopt the commonly used setting $\tau=0.1$. To further investigate its impact, we perform a grid search over $\tau\in{0.5,0.1,0.05,0.01}$. As shown in Tab. 11, $\tau=0.1$ and $\tau=0.05$ yield very similar performance, while both larger and smaller values lead to a noticeable degradation in accuracy.
Table 11: Different values for the temperature parameter. Best in bold.
Temperature $\tau$ Mean loc. (m) Median loc. (m) Mean ori. () Median ori. ()
0.5 9.58 6.18 39.03 15.33
0.1 (default) 3.86 1.75 9.30 1.93
0.05 3.78 1.80 9.25 2.17
0.01 4.58 2.46 13.70 4.78
BEV-plane matching: Transforming ground images into BEV via a homography typically introduces ray-directional distortions, which degrade correspondence quality when matching these BEV-transformed views directly to the aerial image. To validate this, we construct a bird’s-eye-view (BEV) variant of Loc2 by first transforming the ground-level images into BEV representations and then feeding the transformed images into the same feature extractor and matching head, followed by scale-aware Procrustes alignment using the same loss functions. We adopted the BEV transformation from HC-Net (Wang et al., 2024b) with its default hyperparameters.
As shown in the Tab. 12, the BEV variant yields substantially worse performance. The BEV transformation inevitably introduces distortions to elevated or above-ground structures. More importantly, such structures often occupy a large portion of the input image, restricting matching to a small undistorted region or forcing the model to match features on distorted objects, which ultimately degrades performance (see Fig. 7).
Table 12: Comparison between the proposed Loc2 and its BEV variant. Best in bold.
BEV-plane matching 9.49 8.20 53.42 31.84
Homography pose estimation 14.46 13.43 90.98 91.17
Ours 3.86 1.75 9.30 1.93
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 7: Matching results of the BEV variant of our method on the VIGOR same-area test set under unknown orientation. We visualize the top 50 correspondences, ranked by matching score.
Homography-based pose estimation: We next replace our pose estimation module with a homography-based formulation instead of a 2D similarity transformation. Specifically, we use the sampled correspondences to estimate a homography matrix and supervise it through translation and orientation, following a strategy similar to HC-Net (Wang et al., 2024b).
As shown in the Tab. 12, the homography-based formulation leads to a significant drop in localization accuracy. A homography has eight degrees of freedom and, beyond translation and rotation, can model anisotropic scaling and shear. As a result, it can disregard the contour and shape information encoded in the (relative) depth prior. Meanwhile, it also reduces interpretability: the inferred ground layout may become skewed or non-uniformly scaled across different directions in order to satisfy the point correspondences in the aerial view.
@@ Appendix B Test on Non-Road Facing Samples on KITTI
Unknown Cross-area (non-road) CCVPE 8.33 5.99 89.78 93.97
@@ Appendix C Additional Local Feature Matching Results
Failure cases: Typical failure cases are shown in Fig. 10. When there is dense vegetation (Fig. 10 (a),(b)), the matchable information across views is very limited, making precise localization extremely challenging, even for humans. Interestingly, in Fig. 10 (b), our method still correctly matches the gap between two buildings in the ground view to the corresponding gap in the aerial view. In Fig. 10 (c), the two narrow roads with trees and a zebra crossing in front appear visually similar. At the true location, a tree partially blocks the zebra crossing in the aerial view, making it harder to distinguish. As a result, our method matches the zebra crossing to an incorrect one, leading to an erroneous location estimate. In Fig. 10 (d), it is nearly impossible to determine the exact location over the water due to the lack of distinctive features. Finally, Fig. 10(e)–(h) shows that occlusions from dense high-rise buildings make cross-view matching more difficult. Importantly, we believe these scenarios represent common challenges for most cross-view localization methods, rather than limitations specific to our approach. Furthermore, our local feature matching framework provides strong interpretability, enabling straightforward identification of such failure cases and supporting RANSAC-based outlier filtering.
@@
Refer to caption(e)
Refer to caption(f)
Refer to caption(g)
Refer to caption(h)
@@
Appendix D Additional results on cross-dataset generalization.
In addition to Sec. 4.5, we present more qualitative results on cross-dataset generalization. As shown in Fig. 11, our model can establish reliable correspondences on buildings (a), road markings (b, c, g, f), and trees (d, e). The layout alignments also demonstrate our superior localization quality with recovered rotation, translation and scale.
Refer to caption(a)
Refer to caption(b)
Refer to caption(a)
Refer to caption(c)
Refer to caption(b)
Refer to caption(d)
Refer to caption(c)
Refer to caption(d)
Refer to caption(e)
Refer to caption(f)
Refer to caption(e)
Refer to caption(g)
Refer to caption(f)
Refer to caption(h)
Refer to caption(g)
Refer to caption(h)
Figure 11: Direct generalization to the CVACT dataset. We visualize the top 50 correspondences, ranked by matching score, and overlay the ground layout on the aerial image after using the predicted rotation, translation, and scale transformations.
Appendix E Additional Results on Using Relative Depth in Inference
@@
Appendix F Training and Testing with Relative Depth
@@
Appendix G Additional Proof for Scale-Aware Procrustes Alignment
@@
Appendix H Additional Details on InfoNCE Losses
@@
Appendix I Discussion of standard and true digital orthophoto maps
A standard Digital Orthophoto Map (DOM) is an aerial or satellite image that has been orthorectified using a Digital Elevation Model (DEM), which typically represents only the bare-earth terrain. Unlike raw aerial photographs, a DOM has a uniform scale and corrected geometric distortions, enabling it to serve as a metrically accurate base map for overlaying additional geospatial information. The aerial imagery in datasets such as VIGOR (Zhu et al., 2021) and KITTI (Shi and Li, 2022) falls into this category, as these images provide uniform scale, they still contain building facades and occlusions due to the absence of above-ground structures in the underlying DEM.
A True Digital Orthophoto Map (TDOM), in contrast, is generated using a Digital Surface Model (DSM) that captures both the terrain and above-ground objects such as buildings and vegetation. As a result, a TDOM removes facade distortions and yields an image that is geometrically much closer to a true nadir (top-down) projection. Our formulation estimates a similarity transformation (rotation, translation, and scale) between ground-level and aerial points, implicitly assuming that the aerial image behaves like a TDOM.
Despite this approximation, our method mitigates the resulting geometric inconsistencies through two key mechanisms: (1) Deep feature representations, which encode not only local appearance but also contextual information aggregated over large receptive fields, making them less sensitive to small misalignments between aerial and ground views; and (2) Confidence-based correspondence sampling, applied during both training and inference, which ensures that pose estimation relies primarily on high-quality matches. As illustrated in the main paper Fig. 3(c), the model avoids unreliable points on building facades and instead focuses on stable cues such as road markings. Likewise, in examples (a) and (d), streetlights are matched via the pole base to their aerial footprint rather than the poles themselves, which is more susceptible to perspective distortion.
Appendix J Runtime and Memory Usage
@@ Appendix I Discussion of standard and true digital orthophoto maps
Runtime: On a single H100 GPU, our method achieves 14.74 FPS without RANSAC (Fischler and Bolles, 1981) (0.03s for our model and 0.04s for the depth predictor Unik3D (Piccinelli et al., 2025)) and 0.42 FPS with RANSAC. This is faster than the previous local feature matching method FG2 (Xia and Alahi, 2025), which runs at 9.26 FPS without RANSAC and 0.32 FPS with it. Notably, when prioritizing speed, omitting RANSAC results in only a small increase in localization error, e.g., on VIGOR same-area test set with known orientation the mean error increases from 3.06 m to 3.29 m. Furthermore, one can use a faster depth predictor to further reduce inference time.
@@
Appendix K LLM Usage Statement
We used a large language model (LLM) exclusively for text polishing, limited to minor grammar corrections, wording refinements, and improvements in readability. The LLM was not involved in the design of the method, the implementation of experiments, the analysis of results, or the generation of any scientific content. All technical ideas, experiments, and conclusions presented in this work are entirely the authors’ own. The role of the LLM was purely editorial, comparable to language editing support.